OpenStudy (anonymous):

solve for x: 9^2x=27

jimthompson5910 (jim_thompson5910):

One way to do this is to get each base equal to the same number (in this case, we can get them both equal to 3 since 9 = 3^2 and 27 = 3^3) \[\Large 9^{2x} = 27\] \[\Large (3^2)^{2x} = 3^3\] \[\Large 3^{2*2x} = 3^3\] Do you see what to do from here?

OpenStudy (anonymous):

not really

jimthompson5910 (jim_thompson5910):

The two sides are equal. In order for that to be true, the exponents must be equal (since the bases are equal to 3)

jimthompson5910 (jim_thompson5910):

So that means 2*2x = 3 now solve for x

OpenStudy (anonymous):

oh okay thank you

jimthompson5910 (jim_thompson5910):

you're welcome